OpenStudy (anonymous):

a test isworth 125 points . How many points are needed to obtain a score of 85%?

OpenStudy (anonymous):

wat is 125's 85%?

OpenStudy (anonymous):

is 85%= 85/100?

OpenStudy (anonymous):

125=100

jimthompson5910 (jim_thompson5910):

take 85% of 125 to get... 85% of 125 = (85/100)*125 85% of 125 = 0.85*125 85% of 125 = 106.25 If you round to the nearest point, then you have to round up to guarantee you get at least 85% or more So you have to score 107 out of 125

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

125=100